OpenStudy (anonymous):

The bob of a simple pendulum is imparted a velocity 5m/s when it is at its mean position. To what maximum vertical height will it rise on reaching its extreme position if 60% of its energy is lost in overcoming friction of air? (Take g=10m/s^2)

OpenStudy (amistre64):

potential energy is equal to 60% of the kinetic energy ... mgh = .6/2 mv^2 gh = .6/2 v^2 h = .6/(2g) v^2

OpenStudy (anonymous):

is it not mgh =0.4/2 mv^2 since 60% is lost ?